Page 280 Problem #7-14    Using the following equations, graph the constraints, and solve using the corner point approach.

Let: X1 = number of air conditioners to be produced

X2 = number of fans to be produced

Maximize profit = 25X1 + 15X2

subject to 3X1 + 2X2 <= 240 (wiring)

               2X1 + 1X2 <= 140 (drilling)

                     X1, X2 >= 0
Answer #1

Max Profit 25X1 15X2 3x1+2X2 <= 240 (0, 120) x2 70 2X1 + 1X2 <= 140 x2 140 10 x1,X20 12 13 14 15 16 17 18 19 Coordinates (40,Max Profit 25X1 15X2 3X1 2X2 240 70 2x1+ 1x2 <= 140 140 10 x1, X20 13 14 15 16 17 18 19 20 Coordinates Profit -25 B16+15 C16